Find the integral surface of the partial differential equation:(x–y) y^2 p+(y–x)x^2 q=(x^2 +y^2)z through the curve xz=a^2, y=0
1
Expert's answer
2020-02-11T09:25:43-0500

"dx\/((x-y)*y^2)=dy\/((y-x)*x^2)=dz\/((x^2+y^2)*z)"

"dx\/((x-y)*y^2)=dy\/((y-x)*x^2)"

"dy\/dx=-x^2\/y^2"

"dy*y^2=dx*(-x^2)"

"C1=y^3+x^3"

"dx\/((x-y)*y^2)=dz\/((x^2+y^2)*z)"

"dz\/z=dx(x^2+y^2)\/((x-y)*y^2)"

"lnz=1\/y^2*F(x,y)+C2"

y can't be 0 in this equation that is why there is no intersection with curve xz=a^2, y=0.

Answer: no solution.